the equation of a line yhat is perpendicular to the line y=-3x+4
Answer by markburke(3)

The line is in the form y = m times x plus b. b is the y-intercept, the value on the y-axis where the line crosses that y-axis. m is the slope, also is the number in front of the x term.
Parallel lines will have the same slope as the line above. This means they will always have the same m value as the line above. There are a many different lines that can be "parallel", that is drawn with same slope, also described as drawn in the same direction as the line above.) They will have the same "m" as above line, but possible different values of b. Perpendicular lines are 90 degrees rotated from the parallel ones, which means that their slope is just the minus sign version of the parallel lines' slope (opposite sign slope). Think of another example y=2x+1. The perpendicular line to this is y= -2 x + 1, or another perpendicular line is y = -2 x+ 2, and y = -2 x + 3. These are all perpendicular lines to the y = 2 x + 1 line. All of the perpendicular lines have -2 slope, and the original line has 2 slope. The line is in the form y=mx + b, where b is the y-intercept, or the point crossing the y-axis.
So the answer is, an example perpendicular line for the line y = -3 x + 4 , is y = 3 x + 4. Also many answers exist, basically any line which is y = 3 x + (any b value, you pick) would be a perpendicular line.
Notice that the parallel lines (same slope as line) and the perpendicular lines (negative sign slope as the line) are always always 90 degrees rotated on the graph, from each other.